# Ophiuchus: Scalable Modeling of Protein Structures through Hierarchical Coarse-graining SO(3)-Equivariant Autoencoders

### v1

@@ Appendix C Transpose Sequence Convolution
Windows generated through Transpose Convolutions predict anchoring positions and representations that intersect when mapped back to sequence. We resolve those superpositions by taking the mean anchor and mean representations across intersections.
Appendix D Loss Details
Our loss function minimizes differences of internal vector maps $||V(\mathbf{P})-V(\hat{\mathbf{P}})||^{1}_{1}$ between predicted and ground positions, where $V(\mathbf{P})^{i,j}=(\mathbf{P}^{i}-\mathbf{P}^{j})\in\mathbb{R}^{N\times N\times 3}$. Our output algorithm for decoding atoms produces arbitrary symmetry breaks (Appendix A) for permutation invariances of particular side-chain atoms. Because of that, the vector map loss is not directly applicable to the output of our model. Instead, we first compare vector maps internal to individual residues. We consider the alternative permutation of ground truth positions, and pick the solution which minimizes the vector map loss. We then flip ground truth for the rest to apply the loss across residues in all-atom settings. A similar approach to tackling side-chain symmetries is described by [Jumper et al. (2021)].
Appendix E Visualization of Latent Space
Appendix F More on Ablation
@@
In addition to all-atom ablation found in Table 1 we conducted a similar ablation study on models trained on backbone only atoms as shown in Table 2. We found that backbone only models performed slightly better on backbone reconstruction. Furthermore, in Fig. 7.a we compare the relative vector loss between ground truth coordinates and the coordinates reconstructed from the autoencoder with respect to different downsampling factors. We average over different trials and channel rescale factors. As expected, we find that for lower downsampling factors the structure reconstruction accuracy is better. In Fig 7.b we similarly plot the residue recovery accuracy with respect to different downsampling factors. Again, we find the expected result, the residue recovery is better for lower downsampling factors. Notably, the relative change in structure recovery accuracy with respect to different downsampling factors is much lower compared to the relative change in residue recovery accuracy for different downsampling factors. This suggest our model was able to learn much more efficiently a compressed prior for structure as compared to sequence, which coincides with the common knowledge that sequence has high redundancy in biological proteins. In Fig. 7.c we compare the structure reconstrution accuracy across different channel rescaling factors. Interestingly we find that for larger rescaling factors the structure reconstruction accuracy is slightly lower. However, since we trained only for 10 epochs, it is likely that due to the larger number of model parameters when employing a larger rescaling factor it would take somewhat longer to achieve similar results. Finally, in Fig. 7.d we compare the residue recovery accuracy across different rescaling factors. We see that for higher rescaling factors we get a higher residue recovery rate. This suggests that sequence recovery is highly dependant on the number of model parameters and is not easily capturable by efficient structural models.
@@
Appendix G Additional Conformational Interpolation
We provide an additional example from PDBFlex in Fig. 8.
Figure 8: (a) Comparison of chemical Comparison of chemical validity metrics across interpolated structures of Maltose-binding periplasmic protein (PDB ID: 4JKM and PDB 6LF3). We plot results for interpolations on the original data space, on the latent space, and on the autoencoder-reconstructed data space. (b) Direct linear interpolation of coordinates. (c) Decoded linear latent interpolation between the same encoded states as in (b).
Appendix H Diffusion Details
To train an autoencoder for latent diffusion over the miniprotein library, we instantiate an Ophiuchus model with stride $S=3$ and kernel size $K=5$. The model progressively coarsens sequences across lengths [63, 31, 15, 7, 3, 1] while producing representations of dimensionalities [12, 18, 27, 40, 60, 90].
We regularize the learned latent space so that representations are amenable to the range scales of diffusion. Let $\mathbf{V}^{l}_{i}$ denote the $i$-th channel of a vector representation $\mathbf{V}^{l}$. We regularize the autoencoder latent space by optimizing radial and angular components of our vectors:
The first term penalizes vector magnitudes larger than one, and the second term induces vectors to spread angularly. We find these regularizations to significantly help training of the denoising diffusion model. The diffusion model uses Self-Interactions with full tensor product [B].
To compute the scTM scores, we recover 8 sequences using ProteinMPNN for all sampled backbones from our diffusion model. We used a sampling temperature of 0.1 for ProteinMPNN. Unlike the original work, where the sequences where folded using AlphaFold2, we use OmegaFold[Wu et al. (2022)] similar to [Lin & AlQuraishi (2023)].The predicted structures are aligned to the original sampled backbones and TM-Score is calculated for each alignment. For the all-atom generation we calculated the TM-scores by folding the generated sequence using OmegaFold (Fig. 5.b.)

### v2

@@ A.6 Transpose Sequence Convolution
This procedure generates windows of $K$ representations and positions. These windows may intersect in the decoded output. We resolve those superpositions by taking the average position and average representations within intersections.
A.7 Layer Normalization and Residual Connections
Training deep models can be challenging due to vanishing or exploding gradients. We employ layer normalization [Ba et al. (2016)] and residual connections [He et al. (2015)] in order to tackle those challenges. We incorporate layer normalization and residual connections at the end of every Self-Interaction and every convolution. To keep roto-translation equivariance, we use the layer normalization described in Equiformer [Liao & Smidt (2023)], which rescales the $l$ signals independetly within a representation by using the root mean square value of the vectors. We found both residuals and layer norms to be critical in training deep models for large proteins.
A.8 Encoder-Decoder
Below we describe the encoder/decoder algorithm using the building blocks previously introduced.
Input: All-Atom Relative Positions $\mathbf{P}^{\ast}\in\mathbb{R}^{n\times 3}$
Input: Residue Label $\mathbf{R}\in\mathcal{R}$
Output: Latent Representation $(\mathbf{P},\mathbf{V}^{l=0:l_{\textrm{max}}})$
$(\mathbf{P},\mathbf{V}^{0:2})\leftarrow\textrm{All-Atom Encoding}(\mathbf{P}^{\alpha},\mathbf{P}^{\ast},\mathbf{R})$  normal-$\triangleright$ Initial Residue Encoding
$(\mathbf{P},\mathbf{V}^{0:l_{\textrm{max}}})\leftarrow\textrm{Linear}(\mathbf{P},\mathbf{V}^{0:2})$ for $i\leftarrow 1$ to $L$ do
       $(\mathbf{P},\mathbf{V}^{0:l_{\textrm{max}}})\leftarrow\textrm{Self-Interaction}(\mathbf{P},\mathbf{V}^{0:l_{\textrm{max}}})$ $(\mathbf{P},\mathbf{V}^{0:l_{\textrm{max}}})\leftarrow\textrm{Spatial Convolution}(\mathbf{P},\mathbf{V}^{0:l_{\textrm{max}}})$ $(\mathbf{P},\mathbf{V}^{0:l_{\textrm{max}}})\leftarrow\textrm{Sequence Convolution}(\mathbf{P},\mathbf{V}^{0:l_{\textrm{max}}})$
end for
return $\mathbf{(}\mathbf{P},\mathbf{V}^{0:{l_{\max}}})$
Algorithm 7 Encoder
Input: Latent Representation ($\textrm{Latent Representation }(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})$
Output: Protein Structure and Sequence Logits $(\hat{\mathbf{P}}^{\alpha},\hat{\mathbf{P}}^{\ast},\bm{\ell})$
for $i\leftarrow 1$ to $L$ do
       $(\mathbf{P},\mathbf{V}^{0:l_{\textrm{max}}})\leftarrow\textrm{Self-Interaction}(\mathbf{P},\mathbf{V}^{0:l_{\textrm{max}}})$ $(\mathbf{P},\mathbf{V}^{0:l_{\textrm{max}}})\leftarrow\textrm{Spatial Convolution}(\mathbf{P},\mathbf{V}^{0:l_{\textrm{max}}})$ $(\mathbf{P},\mathbf{V}^{0:l_{\textrm{max}}})\leftarrow\textrm{Transpose Sequence Convolution}(\mathbf{P},\mathbf{V}^{0:l_{\textrm{max}}})$
end for
$(\mathbf{P},\mathbf{V}^{0:2})\leftarrow\textrm{Linear}(\mathbf{P},\mathbf{V}^{0:l_{\textrm{max}}})$ $(\hat{\mathbf{P}}^{\alpha},\hat{\mathbf{P}}^{\ast},\bm{\ell})\leftarrow\textrm{All-Atom Decoding}(\mathbf{P},\mathbf{V}^{0:2})$   normal-$\triangleright$ Final Protein Decoding
return $(\hat{\mathbf{P}}^{\alpha},\hat{\mathbf{P}}^{\ast},\bm{\ell})$
Algorithm 8 Decoder
A.9 Variable Sequence Length
We handle proteins of different sequence length by setting a maximum size for the model input. During training, proteins that are larger than the maximum size are cropped, and those that are smaller are padded. The boundary residue is given a special token that labels it as the end of the protein. For inference, we crop the tail of the output after the sequence position where this special token is predicted.
A.10 Time Complexity
We analyse the time complexity of a forward pass through the Ophiuchus autoencoder. Let the list $(\mathbf{P},\mathbf{V}^{0:l_{\textrm{max}}})_{N}$ be an arbitrary latent state with $N$ positions and $N$ geometric representations of dimensionality $D$, where for each $d\in[0,D]$ there are geometric features of degree $l\in[0,l_{\textrm{max}}]$. Note that size of a geometric representation grows as $O(D\cdot l^{2}_{\textrm{max}})$ in memory.
  • The cost of an SO(3)-equivariant linear layer (Fig. 7.b) is $O(D^{2}\cdot l_{\textrm{max}}^{3})$.
  • The cost of a channel-wise tensor square operation (Fig. 7.c) is $O(D\cdot l^{4}_{\textrm{max}})$.
  • In Self-Interaction (Alg. 3), we use a tensor square and project it using a linear layer. The time complexity is given by $O\left(N\cdot(D^{2}\cdot l_{\textrm{max}}^{3}+D\cdot l_{\textrm{max}}^{4})\right)$ for the whole protein.
  • In Spatial Convolution (Alg. 5), a node aggregates geometric messages from $k$ of its neighbors. Resolving the $k$ nearest neighbors can be efficiently done in $O\left((N+k)\log N\right)$ through the k-d tree data structure [Bentley (1975)]. For each residue, its $k$ neighbors prepare messages through linear layers, at total cost $O(N\cdot k\cdot D^{2}\cdot l^{3}_{\textrm{max}})$.
  • In a Sequence Convolution (Alg. 4), a kernel stacks $K$ geometric representations of dimensionality $D$ and linearly maps them to a new feature of dimensionality $\rho\cdot D$, where $\rho$ is a rescaling factor, yielding $O((K\cdot D)\cdot(\rho\cdot D)\cdot l^{3}_{\textrm{max}})=O(K\cdot\rho\cdot D^{2}\cdot l^{3}_{\textrm{max}})$. With length $N$ and stride $S$, the total cost is $O\left(\frac{N}{S}\cdot K\cdot\rho\cdot D^{2}\cdot l^{3}_{\textrm{max}}\right)$.
  • The cost of an Ophiuchus Block is the sum of the terms above,
    $O\left(D\cdot l_{\textrm{max}}^{3}\cdot N\cdot(\frac{K}{S}\cdot\rho D+kD)\right)$
  • An Autoencoder (Alg. 7,8) that uses stride $S$ convolutions for coarsening uses $L=O\left(\log_{S}(N)\right)$ layers to reduce a protein of size $N$ into a single representation. At depth $i$, the dimensionality is given by $D_{i}=\rho^{i}D$ and the sequence length is given by $N_{i}=\frac{N}{S^{i}}$. The time complexity of our Autoencoder is given by geometric sum:
    $O\left(\sum_{i}^{\log_{S}(N)}\left(l_{\textrm{max}}^{3}\rho^{i}D\frac{N}{S^{i}}(K\rho^{i+1}D+k\rho^{i}D)\right)\right)=O\left(Nl_{\textrm{max}}^{3}(K\rho+k)D^{2}\sum_{i}^{\log_{S}(N)}\left(\frac{\rho^{2}}{S}\right)^{i}\right)$
    We are interested in the dependence on the length $N$ of a protein, therefore, we keep only relevant parameters. Summing the geometric series and using the identity $x^{\log_{b}(a)}=a^{\log_{b}(x)}$ we get:
    $=O\left(N\frac{(\frac{\rho^{2}}{S})^{\log_{S}(N)+1}-1}{\frac{\rho^{2}}{S}-1}\right)=\begin{cases}O\left(N^{1+\log_{S}\rho^{2}}\right)&\text{for }\rho^{2}/S>1\\ O\left(N\log_{S}N\right)&\text{for }\rho^{2}/S=1\\ O\left(N\right)&\text{for }\rho^{2}/S<1\end{cases}$
    In most of our experiments we operate in the $\rho^{2}/S<1$ regime.
Appendix B Loss Details
B.1 Details on Vector Map Loss
The Huber Loss [Huber (1992)] behaves linearly for large inputs, and quadratically for small ones. It is defined as:
$\text{HuberLoss}(y,f(x))=\begin{cases}\frac{1}{2}(y-f(x))^{2}&\text{if }|y-f(x)|\leq\delta,\\ \delta\cdot|y-f(x)|-\frac{1}{2}\delta^{2}&\text{otherwise.}\end{cases}$
We found it to significantly improve training stability for large models compared to mean squared error. We use $\delta=0.5$ for all our experiments.
The vector map loss measures differences of internal vector maps $V(\mathbf{P})-V(\hat{\mathbf{P}})$ between predicted and ground positions, where $V(\mathbf{P})^{i,j}=(\mathbf{P}^{i}-\mathbf{P}^{j})\in\mathbb{R}^{\times 3}$. Our output algorithm for decoding atoms produces arbitrary symmetry breaks (Appendix A.2) for positions $\mathbf{P}_{v}^{*}$ and $\mathbf{P}_{u}^{*}$ of atoms that are not orderable. Because of that, a loss on the vector map is not directly applicable to the output of our model, since the order of the model output might differ from the order of the ground truth data. To solve that, we consider both possible orderings of permutable atoms, and choose the one that minimizes the loss. Solving for the optimal ordering is not feasible for the system as a whole, since the number of permutations to be considered scales exponentially with $N$. Instead, we first compute a vector map loss internal to each residue. We consider the alternative order of permutable atoms, and choose the candidate that minimizes this local loss. This ordering is used for the rest of our losses.
B.2 Chemical Losses
We consider bonds, interbond angles, dihedral angles and steric clashes when computing a loss for chemical validity. Let $\mathbf{P}\in\mathbb{R}^{N_{\textrm{Atom}}\times 3}$ be the list of atom positions in ground truth data. We denote $\hat{\mathbf{P}}\in\mathbb{R}^{N_{\textrm{Atom}}\times 3}$ as the list of atom positions predicted by the model. For each chemical interaction, we precompute indices of atoms that perform the interaction. For example, for bonds we precompute a list of pairs of atoms that are bonded according to the chemical profile of each residue. Our chemical losses then take form:
$\mathcal{L}_{\textrm{Bonds}}=\frac{1}{|\mathcal{B}|}\sum_{(v,u)\in\mathcal{B}}\Big{|}\Big{|}\;\;||\hat{\mathbf{P}}_{v}-\hat{\mathbf{P}}_{u}||_{2}-||\mathbf{P}_{v}-\mathbf{P}_{u}||_{2}\;\;\Big{|}\Big{|}^{2}_{2}$
where $\mathcal{B}$ is a list of pair of indices of atoms that form a bond. We compare the distance between bonded atoms in prediction and ground truth data.
$\mathcal{L}_{\textrm{Angles}}=\frac{1}{|\mathcal{A}|}\sum_{(v,u,p)\in\mathcal{A}}||\alpha(\hat{\mathbf{P}}_{v},\hat{\mathbf{P}}_{u}\hat{\mathbf{P}}_{p})-\alpha\left(\mathbf{P}_{v},\mathbf{P}_{u},\mathbf{P}_{p}\right)||^{2}_{2}$
where $\mathcal{A}$ is a list of 3-tuples of indices of atoms that are connected through bonds. The tuple takes the form $(v,u,p)$ where $u$ is connected to $v$ and to $p$. Here, the function $\alpha(\cdot,\cdot,\cdot)$ measures the angle in radians between positions of atoms that are connected through bonds.
$\mathcal{L}_{\textrm{Dihedrals}}=\frac{1}{|\mathcal{D}|}\sum_{(v,u,p,q)\in\mathcal{D}}||\tau(\hat{\mathbf{P}}_{v},\hat{\mathbf{P}}_{u},\hat{\mathbf{P}}_{p},\hat{\mathbf{P}}_{q})-\tau(\mathbf{P}_{v},\mathbf{P}_{u},\mathbf{P}_{p},\mathbf{P}_{q})||^{2}_{2}$
where $\mathcal{D}$ is a list of 4-tuples of indices of atoms that are connected by bonds, that is, $(v,u,p,q)$ where $(v,u)$, $(u,p)$ and $(p,q)$ are connected by bonds. Here, the function $\tau(\cdot,\cdot,\cdot,\cdot)$ measures the dihedral angle in radians.
$\mathcal{L}_{\textrm{Clashes}}=\frac{1}{|\mathcal{C}|}\sum_{(v,u)\in\mathcal{C}}H(r_{v}+r_{u}-||\hat{\mathbf{P}}_{v}-\hat{\mathbf{P}}_{u}||_{2})$
where $H$ is a smooth and differentiable Heaviside-like step function, $\mathcal{C}$ is a list of pair of indices of atoms that are not bonded, and ($r_{v},r_{u}$) are the van der Waals radii of atoms $v$, $u$.
B.3 Regularization
When training autoencoder models for latent diffusion, we regularize the learned latent space so that representations are amenable to the relevant range scales of the source distribution $\mathcal{N}(0,1)$. Let $\mathbf{V}^{l}_{i}$ denote the $i$-th channel of a vector representation $\mathbf{V}^{l}$. We regularize the autoencoder latent space by optimizing radial and angular components of our vectors:
The first term penalizes vector magnitudes larger than one, and the second term induces vectors to spread angularly. We find these regularizations to significantly help training of the denoising diffusion model.
B.4 Total Loss
We weight the different losses in our pipeline. For the standard training of the autoencoder, we use weights:
$\mathcal{L}=10\cdot\mathcal{L}_{\textrm{VectorMap}}+\mathcal{L}_{\textrm{CrossEntropy}}+0.1\cdot\mathcal{L}_{\textrm{Bonds}}+0.1\cdot L_{\textrm{Angles}}+0.1\cdot\mathcal{L}_{\textrm{Dihedrals}}+10\cdot\mathcal{L}_{\textrm{Clashes}}$
For fine-tuning the model, we increase the weight of chemical losses significantly:
$\mathcal{L}=10\cdot\mathcal{L}_{\textrm{VectorMap}}+\mathcal{L}_{\textrm{CrossEntropy}}+1.0\cdot\mathcal{L}_{\textrm{Bonds}}+1.0\cdot L_{\textrm{Angles}}+1.0\cdot\mathcal{L}_{\textrm{Dihedrals}}+100\cdot\mathcal{L}_{\textrm{Clashes}}$
We find that high weight values for chemical losses at early training may hurt the model convergence, in particular for models that operate on large lengths.
Appendix C Details on Autoencoder Comparison
We implement the comparison EGNN model in its original form, following [Satorras et al. (2022)]. We use kernel size $K=3$ and stride $S=2$ for downsampling and upsampling, and follow the procedure described in [Fu et al. (2023)]. For this comparison, we train the models to minimize the residue label cross entropy, and the vector map loss.
Ophiuchus significantly outperforms the EGNN-based architecture. The standard EGNN is SO(3)-equivariant with respect to its positions, however it models features with SO(3)-invariant representations. As part of its message passing, EGNN uses relative vectors between nodes to update positions. However, when downsampling positions, the total number of relative vectors available reduces quadratically, making it increasingly challenging to recover coordinates. Our method instead uses SO(3)-equivariant feature representations, and is able to keep 3D vector information in features as it coarsens positions. Thus, with very few parameters our model is able to encode and recover protein structures.
Appendix D More on Ablation
@@
In addition to all-atom ablation found in Table 2 we conducted a similar ablation study on models trained on backbone only atoms as shown in Table 4. We found that backbone only models performed slightly better on backbone reconstruction. Furthermore, in Fig. 9.a we compare the relative vector loss between ground truth coordinates and the coordinates reconstructed from the autoencoder with respect to different downsampling factors. We average over different trials and channel rescale factors. As expected, we find that for lower downsampling factors the structure reconstruction accuracy is better. In Fig 9.b we similarly plot the residue recovery accuracy with respect to different downsampling factors. Again, we find the expected result, the residue recovery is better for lower downsampling factors.
Notably, the relative change in structure recovery accuracy with respect to different downsampling factors is much lower compared to the relative change in residue recovery accuracy for different downsampling factors. This suggest our model was able to learn much more efficiently a compressed prior for structure as compared to sequence, which coincides with the common knowledge that sequence has high redundancy in biological proteins. In Fig. 9.c we compare the structure reconstrution accuracy across different channel rescaling factors. Interestingly we find that for larger rescaling factors the structure reconstruction accuracy is slightly lower.
However, since we trained only for 10 epochs, it is likely that due to the larger number of model parameters when employing a larger rescaling factor it would take somewhat longer to achieve similar results. Finally, in Fig. 9.d we compare the residue recovery accuracy across different rescaling factors. We see that for higher rescaling factors we get a higher residue recovery rate. This suggests that sequence recovery is highly dependant on the number of model parameters and is not easily capturable by efficient structural models.
@@
Appendix E Latent Space Analysis
E.1 Visualization of Latent Space
Figure 11: Latent Conformational Interpolation. Top: Comparison of chemical validity metrics across interpolated structures of Nuclear Factor of Activated T-Cell (NFAT) Protein (PDB ID: 1S9K and PDB ID: 2O93). Bottom: Comparison of chemical validity metrics across interpolated structures of Maltose-binding periplasmic protein (PDB ID: 4JKM and PDB 6LF3). For both proteins, we plot results for interpolations on the original data space, on the latent space, and on the autoencoder-reconstructed data space.
Appendix F Latent Diffusion Details
We train all Ophiuchus diffusion models with learning rate $\textrm{lr}=1\times 10^{-3}$ for 10,000 epochs. For denoising networks we use Self-Interactions with the chunked-channel tensor square operation (Alg. 3).
Our tested models are trained on two different datasets. The MiniProtein scaffolds dataset consists of 66k all-atom structures of sequence length between 50 and 65, and composes of diverse folds and sequences across 5 secondary structure topologies, and is introduced by [Cao et al. (2022)]. We also train a model on the data curated by [Yim et al. (2023)], which consists of approximately 22k proteins, to compare Ophiuchus to the performance of FrameDiff and RFDiffusion in backbone generation for proteins up to 485 residues.
F.1 Self-Consistency Scores
To compute the scTM scores, we recover 8 sequences using ProteinMPNN for 500 sampled backbones from the tested diffusion models. We used a sampling temperature of 0.1 for ProteinMPNN. Unlike the original work, where the sequences where folded using AlphaFold2, we use OmegaFold [Wu et al. (2022b)] similar to [Lin & AlQuraishi (2023)]. The predicted structures are aligned to the original sampled backbones and TM-Score and RMSD is calculated for each alignment. To calculate the diversity measurement, we hierarchically clustered samples using MaxCluster. Diversity is defined as the number of clusters divided by the total number of samples, as described in [Yim et al. (2023)].
F.2 MiniProtein Model
In Figure (13) we show generated samples from our miniprotein model, and compare the marginal distribution of our predictions and ground truth data. In Figure (14.b) we show the distribution of TM-scores for joint sampling of sequence and all-atom structure by the diffusion model. We produce marginal distributions of generated samples Fig.(14.e) and find them to successfully approximate the densities of ground truth data. To test the robustness of joint sampling of structure and sequence, we compute self-consistency TM-Scores[Trippe et al. (2023)]. 54% of our sampled backbones have scTM scores > 0.5 compared to 77.6% of samples from RFDiffusion. We also sample 100 proteins between 50-65 amino acids in 0.21s compared to 11s taken by RFDiffusion on a RTX A6000.
F.3 Backbone Model
We include metrics on designability in Figure 12.
Figure 12: Designability of Sampled Backbones . (a) To analyze the designability of our sampled backbones we show a plot of scTM vs plDDT. (b) To analyze the composition of secondary structures in the samples we show a plot of helix percentage in a sample vs beta-sheet percentage
[…]
Appendix G Visualization of All-Atom Reconstructions
Figure 15: Reconstruction of 128-length all-atom proteins. Model used for visualization reconstructs all-atom protein structures from coarse representations of 120 scalars and 120 3D vectors.
Appendix H Visualization of Backbone Reconstructions
Figure 16: Reconstruction of 485-length protein backbones. Model used for visualization reconstructs large backbones from coarse representations of 160 scalars and 160 3D vectors.
Appendix I Visualization of Random Backbone Samples
Figure 17: Random Samples from an Ophiuchus 485-length Backbone Model. Model used for visualization samples SO(3) representations of 160 scalars and 160 3D vectors. We measure 0.46 seconds to sample a protein backbone.